Fermi National Laboratory


Accelerator Update for February 18, 2005 - February 21, 2005

Friday February 18
The day shift began with Operations monitoring store 3987 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with MiniBooNE and the Meson MCenter and MTest experiments taking beam.

At 10:56 AM, Operators reported that they had noticed some Booster beam irregularities. They traced the problem back to the H- Source. Experts will investigate further during the next shot setup.

At 11:19 AM, the cryo system technicians at CHL reported that a belt had failed on the Tevatron (TeV) sector C2 wet engine. Before operators could respond, the TeV quenched at C2.

Operations received permission to operate the NuMI beamline at 12:03 PM. Unfortunately, a power supply problem with HV101 kept them from taking beam.

At 1:13 PM, Operations reported the loss of the stash due to a beam valve closing. Experts said that an access would be needed to make repairs.

Linac experts accessed the H- Source at 2:06 PM. Booster experts also took the time to work on a RF power supply (BRF4) and a Debuncher.

Operations put the TeV into shot setup at 2:39 PM.

Operations established store 3989 at 4:58 PM with an initial luminosity of 72.87E30. Pbar resumed stacking.

Operations reported at 5:58 PM that NuMI experts began commissioning their beamline.

At 9:15 PM, Operations reported that a Linac RF station (KRF5) tripped off due to a spark. It tripped a few more times at 9:23 PM and then again at 9:44 PM.

Operations reported at 10:20 PM that they had established a multibatch extraction to NuMI. Experts began a study period.

Saturday February 19
The midnight shift began with Operations monitoring store 3989, with Pbar stacking, with Recycler and NuMI experts conducting studies, and with MiniBooNE and MCenter taking beam.

At 1:38 AM, KRF5 tripped off. It reset okay.

At 7:35 AM, KRF4 tripped off. It reset okay.

At 8:39 AM, LRF4 tripped off. It reset okay.

At 9:44 AM, KRF6 tripped off twice. It reset okay.

At 9:55 AM, KRF6 tripped off for the first of five times in an hour and a half. It reset okay each time.

At 11:08 AM, KRF4 tripped off. Operators eventually reset the station.

At 11:38 AM, KRF6 tripped off and wouldn't reset. Operations contacted experts.

Experts began their access into the Main Injector (MI) and to the Pbar AP1 transfer line at 1 PM.

The MI and AP1 accesses ended and Operations began turning on MI at 2:30 PM.

Operations established beam to MiniBooNE at 2:44 PM.

Pbar resumed stacking at 3:41 PM.

At 4:15 PM, Operations called a TeV expert about an emittance problem that may have been due to the dampers.

At 4:20 PM, Operations turned off the TeV dampers.

Operations resumed sending beam to Switchyard (SY) at 4:25 PM.

Operations established a proton stash into the Recycler at 4:35 PM for study purposes.

At 6:50 PM, a Booster vacuum burst tripped off BRF7 & 8. Operators managed to reset the stations.

The Run Coordinator gave a Pbar expert permission to conduct a stacking study at 7:05 PM.

Operations terminated store 3989 at 7:09 PM. TeV experts began a study period.

At 8:38 PM, operators reported a LCW leak on a Linac RF station (KRF6). An expert was called in.

Operations put the TeV through a dry squeeze at 9:24 PM.

Operations put the TeV into shot setup at 9:53 PM.

Operations established store 3991 at 11:57 PM with an initial luminosity of 72.27E30.

Sunday February 20
The midnight shift began with Operations monitoring store 3991, with a Pbar expert conducting stacking studies, and with Linac off due to a KRF6 LCW leak.

At 1:07 AM, the TeV suffered a sector A1 quench. After an investigation, operators determined that the quench was due to a kicker (C:AAK1) prefire.

The Linac expert reported at 1:33 AM that he had finished work on KRF6. Operations established beam to Pbar, MiniBooNE, and to SY.

Operations resumed sending beam to NuMI at 2:11 AM.

At 3:19 AM KRF5 tripped off. It reset okay. It tripped off again at 4:48 AM. It reset okay.

TeV damper studies began around 7 AM.

Operations reported at 8:40 AM that they had loaded a 36x0 study-store into the TeV.

Operations terminated the study-store at 10:12 and then established another.

Operations terminated the study-store 12:18 PM and then put the TeV through a dry squeeze.

Operations put the TeV into shot setup at 12:56 PM. (Shot setup disables beam to SY and NuMI.)

Operations established store 3994 at 4:39 PM with an initial luminosity of 64.18E30. Pbar resumed stacking.

Operations established beam to NuMI and MCenter at 4:55 PM.

At 5:29 PM, the TeV suffered at sector D1 quench. Operators believe that the quench may have been caused by an A49H separator polarity switch.

Two TeV experts disconnected the polarity switch for the A49H separator at 8:56 PM.

Operations put the TeV into a dry squeeze at 9:15 PM and then into a wet squeeze at 10:38 PM.

Monday February 21
The midnight shift began with TeV checkout, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations put the TeV into shot setup at 1:03 AM.

Operations established store 3997 at 3:13 AM with an initial luminosity of 58.83E30. Pbar resumed stacking.

Both NuMI and MCenter resumed taking by 3:29 AM.

NuMI experts said that they were done with beam at 7:48 AM.

At 8 AM, all the Linac Klystron stations tripped off due to a water skid trip. Beam resumed at 8:15 AM.

The Plan for Monday
The plan for today is to stash antiprotons into the Recycler, to monitor the store, to stack, and to allow NuMI study time. After NuMI completes their study, Operations will put the TeV into shot setup.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 2/21/2005   email Fermilab